OpenStudy (annalee607):

Find the scale factor of a prism with the surface area of 81m(squared) to a similar prism with the surface area of 361m(squared)

OpenStudy (qwertty123):

sqrt(361) : sqrt(81)

OpenStudy (annalee607):

How do I solve that

OpenStudy (annalee607):

Nevermind thank you. Will you help me with another?

OpenStudy (qwertty123):

9:19 given the smaller surface area is mentioned 1st.

OpenStudy (annalee607):

Will you help me with another question please

OpenStudy (qwertty123):

^.^ sure

OpenStudy (annalee607):

Are the two figures similar? If so,give the similarity ratio of the smaller figure to the larger figure.

OpenStudy (mathstudent55):

To be similar, all corresponding lengths must be in the same ratio.

OpenStudy (qwertty123):

Welp he just said what I was going to say...

OpenStudy (annalee607):

So it wouldn't be corresponding?

OpenStudy (mathstudent55):

The large figure has length 34. The small figure has length 17. The ratio of lengths is 34/17 = 2/1 Now check if the ratios of the widths and heights are also the same.

OpenStudy (qwertty123):

Is 17/34 equal to 7/42, also is it equal to 2/6?

OpenStudy (mathstudent55):

Now let's look at heights. The height of the larger solid is 42. The height of the smaller solid is 7. The ratio is 42/7 = 6/1 Since the ratio 6/1 of the heights is not the same as the ratio 2/1 of the lengths, the solids are not similar. We don't even need to check the ratio of the depths.

OpenStudy (annalee607):

Thank you what about with this one?

OpenStudy (qwertty123):

WHAT WAS YOUR ANSWER FOR THE FIST ONE?

OpenStudy (qwertty123):

ahh caps.. whoops :/

OpenStudy (annalee607):

No they are not similar so there is no ratio

OpenStudy (qwertty123):

bingo!

OpenStudy (annalee607):

Thank you!CAn you help with this one?

OpenStudy (mathstudent55):

Look at both cylinders. Each cylinder has two given dimensions. Can you name two dimensions that are corresponding, one in the large cylinder, and one in the small cylinder?

OpenStudy (qwertty123):

Yes ^

OpenStudy (annalee607):

4:6.4

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Correct. The heights are 4 and 6.4, and they are corresponding dimensions. Now can you name the other two corresponding dimensions?

OpenStudy (qwertty123):

You are correct! :)

OpenStudy (mathstudent55):

h/H = 4/6.4 Now what is r/R = ?

OpenStudy (annalee607):

2:3.2

OpenStudy (mathstudent55):

Correct. If the cylinders are similar, then both ratios must be equal. Is 4/6.4 equal to 2/3.2?

OpenStudy (annalee607):

no

OpenStudy (mathstudent55):

\(\dfrac{2}{3.2} = \dfrac{2}{3.2} \times \dfrac{2}{2} = \dfrac{4}{6.4} \) Since multiplying the fraction 2/3.2 equals 4/6.4, then the fractions 2/3.2 and 4/6.4 are equivalent fractions. This means that the ratio of the radii and the ratio of the heights are the same. That makes the cylinders similar.